# Prediction-Powered Adaptive Shrinkage Estimation

### v1

@@ 3 Statistical Guiding Principles and Connections to Prior Work
Table 1: Estimator comparison in the stylized Gaussian model of Section 3.
@@
An important statistical idea is to improve $\hat{\theta}^{\text{cl}}$ via obtaining further information to intercept some of its noise, say $\xi$, and replacing $\hat{\theta}^{\text{cl}}$ by $\hat{\theta}^{\text{cl}}-\xi$ which has MSE $\sigma_{\varepsilon}^{2}$ and remains unbiased for $\theta$. This idea lies at the heart of approaches such as control variates in simulation (Lavenberg and Welch, 1981; Hickernell et al., 2005), variance reduction in randomized controlled experiments via covariate adjustment (Lin, 2013) and by utilizing pre-experiment data (Deng et al., 2013, CUPED), as well as model-assisted estimation in survey sampling (Cochran, 1977; Breidt and Opsomer, 2017). It is also the idea powering PPI and related methods: the unlabeled dataset and the predictive model are used to intercept some of the noise in the classical statistical estimator $\hat{\theta}^{\text{cl}}\triangleq\bar{Y}$; compare to Eq. (2) with $\xi\,\triangleq\,\bar{Z}^{f}-\tilde{Z}^{f}$. We refer to Ji et al. (2025) for an informative discussion of how PPI relates to traditional ideas in semi-parametric inference as in e.g., Robins et al. (1994).
@@
$\mathbb{E}\!\left[\theta\mid\hat{\theta}^{\text{cl}}\right]=\frac{\sigma_{\theta}^{2}}{\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2}+\sigma_{\theta}^{2}}\hat{\theta}^{\text{cl}},$
@@
Now suppose that the variance of the prior, $\sigma_{\theta}^{2}$, is unknown but we observe data from multiple related problems generated from the same model and indexed by $j=1,\dotsc,m$, say, $\theta_{j}\,{\buildrel\text{iid}\over{\sim}\,}\mathcal{N}(0,\sigma_{\theta}^{2})$ and $\hat{\theta}^{\text{cl}}_{j}\,{\buildrel\text{ind}\over{\sim}\,}\mathcal{N}(\theta_{j},\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2})$. Then an EB analysis can mimic the MSE of the oracle Bayesian that has full knowledge of the prior. To wit, we can estimate $\sigma_{\theta}^{2}$ as
$\hat{\sigma}_{\theta}^{2}=\left\{\frac{1}{m-2}\sum_{j=1}^{m}(\hat{\theta}^{\text{cl}}_{j})^{2}\right\}-(\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2}),$
and then consider a plug-in approximation of the Bayes rule, $\hat{\theta}_{j}^{\text{JS}}=\hat{\mathbb{E}}[\theta_{j}\mid\hat{\theta}^{\text{cl}}_{j}]:=\{\hat{\sigma}_{\theta}^{2}/(\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2}+\hat{\sigma}_{\theta}^{2})\}\hat{\theta}^{\text{cl}}$. The resulting estimator is the celebrated James-Stein estimator (James and Stein, 1961; Efron and Morris, 1973), whose risk is very close to the Bayes risk under the hierarchical model. The James-Stein estimator also always dominates the classical estimator under a frequentist evaluation of compound risk in (9) under the assumption that $\hat{\theta}^{\text{cl}}_{j}\,{\buildrel\text{ind}\over{\sim}\,}\mathcal{N}(\theta_{j},\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2})$:
$\mathcal{R}_{m}(\boldsymbol{\hat{\theta}^{\text{JS}}},\boldsymbol{\theta})<\mathcal{R}_{m}(\boldsymbol{\hat{\theta}^{\text{cl}}},\boldsymbol{\theta})\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ for all\leavevmode\nobreak\ \leavevmode\nobreak\ }\boldsymbol{\theta}\in\mathbb{R}^{m}.$
@@
Instead of using the same prior for each problem, we may try to sharpen the prior and increase its relevance (Efron, 2011) by using further information $\phi$. In the stylized example, as seen in Table 1, such an approach reduces the variance of the prior from $\sigma_{\theta}^{2}$ to $\sigma_{\theta\mid\phi}^{2}<\sigma_{\theta}^{2}$ with corresponding MSE reduction of the Bayes estimator. With multiple related problems, such a strategy can be instantiated via EB shrinkage toward an informative but biased predictor (Fay III and Herriot, 1979; Green and Strawderman, 1991; Mukhopadhyay and Maiti, 2004; Kou and Yang, 2017; Rosenman et al., 2023). The strategy of this form that is closest to our proposal is the covariate-powered EB approach of Ignatiadis and Wager (2019). Therein (following the notation of Section 2.2), the analyst has access to classical estimators $\bar{Y}_{j}$, $j\in[m]$, and problem-specific covariates $W_{j}$ and seeks to shrink $\bar{Y}_{j}$ toward ML models that predict $\bar{Y}_{j}$ from $W_{j}$. By contrast, in our setting we have observation-level covariates $X_{ij}$ and the ML model operates on these covariates. In principle one could simultaneously use both types of covariates: problem-specific and observation-specific.
Combined variance reduction (VR) and prior information (P).
One can shrink the variance reduced estimator $\hat{\theta}^{\text{cl}}-\xi$ toward 00 via $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}}-\xi]=\{\sigma_{\theta}^{2}/(\sigma_{\varepsilon}^{2}+\sigma_{\theta}^{2})\}(\hat{\theta}^{\text{cl}}-\xi)$. In the context of PPI, variance reduction and prior information (with a more heavy-tailed prior) are used by Cortinovis and Caron (2025) within the Frequentist-Assisted by Bayes (FAB) framework of Yu and Hoff (2018).  Cortinovis and Caron (2025) only consider a single problem and do not pursue an empirical Bayes approach.
Combined variance reduction (VR), prior (P), and contextual prior information (CP).
Finally, in our stylized example, we can get the smallest MSE (last row of Table 1) by using both variance reduction, shrinkage, and a contextual prior. In that case, the Bayes estimator takes the form,
$\mathbb{E}\!\left[\theta\mid\hat{\theta}^{\text{cl}}-\xi,\phi\right]\,=\,\frac{\sigma_{\theta\mid\phi}^{2}}{\sigma_{\varepsilon}^{2}+\sigma_{\theta\mid\phi}^{2}}(\hat{\theta}^{\text{cl}}-\xi)+\frac{\sigma_{\varepsilon}^{2}}{\sigma_{\varepsilon}^{2}+\sigma_{\theta\mid\phi}^{2}}\mathbb{E}\!\left[\theta\mid\phi\right].$ (12)
EB ideas can be used to mimic the estimator above and provide the starting point for our proposal that we describe next.
@@ 4.1 The Within Problem Power-Tuning Stage
Note that the classical estimator has risk $\sigma_{j}^{2}/n_{j}$ and so is outperformed whenever $\delta_{j}(\lambda)<0$. We can analytically solve for the optimal $\lambda$, which yields
@@
The formulation of the above PT estimators is well understood in the single problem setting (Angelopoulos et al., 2024; Miao et al., 2024a). In PAS, we execute this stage separately for each problem, as the optimal power-tuning parameter is data-dependent and varies case by case.
@@ 4.2 The Across Problem Adaptive Shrinkage Stage
The PT estimator derived in Section 4.1 already possesses many appealing properties: it is unbiased and has lower variance than both the classical estimator and vanilla PPI. However, as our setting involves working with many parallel problems together, there is the opportunity of further MSE reduction by introducing bias in a targeted way. Concretely, based on the PT estimator obtained in Section 4.1, we consider a class of shrinkage estimators
$\hat{\theta}_{j,\omega}^{\mathrm{PAS}}:=\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}+(1-\omega_{j})\tilde{Z}^{f}_{j},\quad\omega_{j}:=\omega_{j}(\omega)=\frac{\omega}{\omega+\tilde{\sigma}^{2}_{j}},\quad\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\omega}:=(\hat{\theta}_{1,\omega}^{\mathrm{PAS}},\ldots,\hat{\theta}_{m,\omega}^{\mathrm{PAS}})^{\intercal},$ (15)
for any $\omega\geq 0$. The motivation is to formally match the form of the Bayes estimator with variance reduction and contextual prior information in (12) with the following (approximate) analogies:
$\hat{\theta}^{\text{cl}}-\xi$ $\longleftrightarrow\hat{\theta}^{\mathrm{PT}}_{j},\;\,$ $\mathbb{E}\!\left[\theta\mid\phi\right]\longleftrightarrow\tilde{Z}^{f}_{j},\;\,$ (16)
@@
    Fixing $\omega$, any problem whose PT estimator has higher variance possesses smaller $\omega_{j}$ and shrinks more towards $\tilde{Z}^{f}_{j}$; a smaller variance increases $\omega_{j}$ and makes the final estimator closer to $\hat{\theta}_{j}^{\mathrm{PT}}$.
@@
    Fixing all the problems, increasing $\omega$ has an overall effect of recovering $\hat{\theta}_{j}^{\mathrm{PT}}$ (full recovery when $\omega\to\infty$), and setting $\omega=0$ recovers $\tilde{Z}^{f}$.
The above establishes the conceptual importance of $\omega$ and if we could choose $\omega$ in an optimal way,
@@
then the resulting estimator $\hat{\boldsymbol{\theta}}_{\omega^{*}}^{\mathrm{PAS}}$ would satisfy all our desiderata. And while the above construction is not feasible as the compound risk function in (9) depends on the unknown $\boldsymbol{\eta},\boldsymbol{\theta}$, we can make progress by pursuing a classical statistical idea: we can develop an unbiased estimate of the compound risk (Mallows, 1973; Stein, 1981; Efron, 2004) and then use that as a surrogate for tuning $\omega$.
@@
$\mathrm{CURE}\left(\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\omega}\right)$ $:=\frac{1}{m}\sum_{j=1}^{m}\left[(2\omega_{j}-1)\tilde{\sigma}^{2}_{j}+2(1-\omega_{j})\tilde{\gamma}_{j}+(1-\omega_{j})^{2}\big{(}\hat{\theta}_{j,\omega_{j}}^{\mathrm{PT}}-\tilde{Z}^{f}_{j}\big{)}^{2}\right],$
where both the formula and our nomenclature (“correlation-aware”) highlight the fact that we must account for the potentially non-zero covariance between shrinkage source $\hat{\theta}^{\mathrm{PT}}_{j}$ and target $\tilde{Z}^{f}_{j}$,
$\tilde{\gamma}_{j}:=\mathrm{Cov}_{\eta_{j}}[\hat{\theta}_{j}^{\mathrm{PT}},\tilde{Z}^{f}_{j}]=\lambda_{j}^{*}\mathrm{Var}_{\eta_{j}}[\tilde{Z}^{f}_{j}]=\frac{\gamma_{j}}{n_{j}+N_{j}}.$ (17)
@@
See Appendices A and B.1 for the proof and motivation. With Theorem 4.1 in hand, we now have a systematic strategy of picking $\omega$ by minimizing CURE, following the paradigm of tuning parameter selection via minimization of an unbiased risk estimate as advocated e.g., by Li (1985); Donoho and Johnstone (1995); Xie et al. (2012); Candès et al. (2013); Ignatiadis and Wager (2019): [^8] The connection to EB is the following. Xie et al. (2012) and Tibshirani and Rosset (2019) explain that James-Stein-type estimators may be derived by tuning $\sigma_{\theta}^{2}$ (in Section 3) via minimization of Stein’s (1981) unbiased risk estimate (SURE).
@@
Even though $\hat{\omega}$ does not admit a closed-form expression, the one-dimensional minimization can be efficiently carried out numerically (e.g. grid search). The final PAS estimator is then given by
@@
0:  $(X_{ij},Y_{ij})_{i=1}^{n_{j}}$, $(\tilde{X}_{ij})_{i=1}^{N_{j}}$, $\rho_{j},\tau_{j},\sigma_{j}$ for $j\in[m]$, predictive model $f$
@@
5:    $\lambda_{j}^{*}=\texttt{{get\_pt\_param}}(\rho_{j},\tau_{j},n_{j},N_{j})$
@@
10:  $\hat{\omega}=\texttt{{get\_shrink\_param}}(\{\hat{\theta}_{j}^{\mathrm{PT}}\}_{j=1}^{m},\{\tilde{Z}^{f}_{j}\}_{j=1}^{m},\{\tilde{\sigma}_{j}^{2}\}_{j=1}^{m}$ )

### v2

@@ 3 Statistical Guiding Principles & Prior Work
Table 1: Estimator comparison in the stylized model of Section 3.
@@
An important statistical idea is to improve $\hat{\theta}^{\text{cl}}$ via obtaining further information to intercept some of its noise, say $\xi$, and replacing $\hat{\theta}^{\text{cl}}$ by $\hat{\theta}^{\text{cl}}-\xi$ which has MSE $\sigma_{\varepsilon}^{2}$ and remains unbiased for $\theta$. This idea lies at the heart of approaches such as control variates in simulation (Lavenberg & Welch, 1981; Hickernell et al., 2005), variance reduction in randomized controlled experiments via covariate adjustment (Lin, 2013) and by utilizing pre-experiment data (Deng et al., 2013, CUPED), as well as model-assisted estimation in survey sampling (Cochran, 1977; Breidt & Opsomer, 2017). It is also the idea powering PPI and related methods: the unlabeled dataset and the predictive model are used to intercept some of the noise in the classical statistical estimator $\hat{\theta}^{\text{cl}}\triangleq\bar{Y}$; compare to Eq. (2) with $\xi\,\triangleq\,\bar{Z}^{f}-\tilde{Z}^{f}$. We refer to Ji et al. (2025) and Gronsbell et al. (2025) for informative discussions of how PPI relates to traditional ideas in semi-parametric inference as in e.g., Robins et al. (1994).
@@
$\mathbb{E}\big{[}\theta\mid\hat{\theta}^{\text{cl}}\big{]}=\frac{\sigma_{\theta}^{2}}{\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2}+\sigma_{\theta}^{2}}\hat{\theta}^{\text{cl}},$
@@
Now suppose that the variance of the prior, $\sigma_{\theta}^{2}$, is unknown but we observe data from multiple related problems generated from the same model and indexed by $j\in[m]$, say, $\theta_{j}\,{\buildrel\text{iid}\over{\sim}\,}\mathcal{N}(0,\sigma_{\theta}^{2})$ and $\hat{\theta}^{\text{cl}}_{j}\,{\buildrel\text{ind}\over{\sim}\,}\mathcal{N}(\theta_{j},\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2})$. Then an EB analysis can mimic the MSE of the oracle Bayesian that has full knowledge of the prior. To wit, we can estimate $\sigma_{\theta}^{2}$ as
$\hat{\sigma}_{\theta}^{2}=\bigg{\{}\frac{1}{m-2}\sum_{j=1}^{m}(\hat{\theta}^{\text{cl}}_{j})^{2}\bigg{\}}-(\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2}),$
and then consider a plug-in approximation of the Bayes rule, $\hat{\theta}_{j}^{\text{JS}}=\hat{\mathbb{E}}[\theta_{j}\mid\hat{\theta}^{\text{cl}}_{j}]:=\{\hat{\sigma}_{\theta}^{2}/(\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2}+\hat{\sigma}_{\theta}^{2})\}\hat{\theta}^{\text{cl}}$. The resulting estimator is the celebrated James-Stein estimator (James & Stein, 1961; Efron & Morris, 1973), whose risk is very close to the Bayes risk under the hierarchical model for large $m$ (Efron, 2010, equation (1.25)). The James-Stein estimator also always dominates the classical estimator under a frequentist evaluation of compound risk in (9) under the assumption that $\hat{\theta}^{\text{cl}}_{j}\,{\buildrel\text{ind}\over{\sim}\,}\mathcal{N}(\theta_{j},\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2})$ and $m\geq 3$:
[…]
@@
Instead of using the same prior for each problem, we may try to sharpen the prior and increase its relevance (Efron, 2011) by using further contextual information $\phi$. In the stylized example, as seen in Table 1, such an approach reduces the variance of the prior from $\sigma_{\theta}^{2}$ to $\sigma_{\theta\mid\phi}^{2}<\sigma_{\theta}^{2}$ with corresponding MSE reduction of the Bayes estimator $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}},\phi]$. With multiple related problems, such a strategy can be instantiated via EB shrinkage toward an informative but biased predictor (Fay III & Herriot, 1979; Green & Strawderman, 1991; Mukhopadhyay & Maiti, 2004; Kou & Yang, 2017; Rosenman et al., 2023). The strategy of this form that is closest to our proposal is the covariate-powered EB approach of Ignatiadis & Wager (2019), recently applied to large language model evaluation by Fogliato et al. (2024). Therein (following the notation of Section 2.2), the analyst has access to classical estimators $\bar{Y}_{j}$, $j\in[m]$, and problem-specific covariates $W_{j}$ and seeks to shrink $\bar{Y}_{j}$ toward ML models that predict $\bar{Y}_{j}$ from $W_{j}$. By contrast, in our setting we have observation-level covariates $X_{ij}$ and the ML model operates on these covariates. In principle one could simultaneously use both types of covariates: problem-specific and observation-specific.
Combine variance reduction (VR) and prior information (P).
One can shrink the variance reduced estimator $\hat{\theta}^{\text{cl}}-\xi$ toward 00 via $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}}-\xi]=\{\sigma_{\theta}^{2}/(\sigma_{\varepsilon}^{2}+\sigma_{\theta}^{2})\}(\hat{\theta}^{\text{cl}}-\xi)$. In the context of PPI, variance reduction and prior information (with a more heavy-tailed prior) are used by Cortinovis & Caron (2025) within the Frequentist-Assisted by Bayes (FAB) framework of Yu & Hoff (2018)Cortinovis & Caron (2025) only consider a single problem and do not pursue an empirical Bayes approach.
Combine P, CP, and VR together.
Finally, in our stylized example, we can get the smallest MSE (last row of Table 1) by using both variance reduction, shrinkage, and a contextual prior. In that case, the Bayes estimator $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}}-\xi,\phi]$ takes the form,
$\frac{\sigma_{\theta\mid\phi}^{2}}{\sigma_{\varepsilon}^{2}+\sigma_{\theta\mid\phi}^{2}}(\hat{\theta}^{\text{cl}}-\xi)+\frac{\sigma_{\varepsilon}^{2}}{\sigma_{\varepsilon}^{2}+\sigma_{\theta\mid\phi}^{2}}\mathbb{E}\!\left[\theta\mid\phi\right].$ (12)
EB ideas can be used to mimic the estimator above and provide the starting point for the proposal we describe next.
@@ 4.1 The Within Problem Power-Tuning Stage
Note that the classical estimator has risk $\sigma_{j}^{2}/n_{j}$ and gets outperformed whenever $\delta_{j}(\lambda)<0$. We can further analytically solve for the optimal $\lambda$, which yields
@@
The formulation of the above PT estimators is well understood in the single problem setting (Angelopoulos et al., 2024; Miao et al., 2024a). In PAS, we execute this stage separately for each problem, as the optimal power-tuning parameter is problem-dependent and varies case by case.
@@ 4.2 The Across Problem Adaptive Shrinkage Stage
The PT estimator derived in Section 4.1 already possesses many appealing properties: it is unbiased and has lower variance than both the classical estimator and vanilla PPI. However, as our setting involves working with many parallel problems together, there is the opportunity of further MSE reduction by introducing bias in a targeted way. [^9] See Section A.2.1 for some explanations about why we can improve MSE by borrowing information across problems. Concretely, based on the PT estimator obtained in Section 4.1, we consider a class of shrinkage estimators $\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\omega}:=(\hat{\theta}_{1,\omega}^{\mathrm{PAS}},\ldots,\hat{\theta}_{m,\omega}^{\mathrm{PAS}})^{\intercal}$, where for any $\omega\geq 0$,
$\hat{\theta}_{j,\omega}^{\mathrm{PAS}}$ $:=\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}+(1-\omega_{j})\tilde{Z}^{f}_{j},$ (15)
$\text{with}\quad\omega_{j}$ $\equiv\omega_{j}(\omega):=\frac{\omega}{\omega+\tilde{\sigma}^{2}_{j}}.$
The motivation is to formally match the form of the Bayes estimator with variance reduction and contextual prior information in (12) with the following (approximate) analogies: [^10] We comment more on these analogies in Section A.2.2.
$\hat{\theta}^{\text{cl}}-\xi$ $\longleftrightarrow\hat{\theta}^{\text{PT}}_{j},\;\,$ $\mathbb{E}\!\left[\theta\mid\phi\right]\longleftrightarrow\tilde{Z}^{f}_{j},\;\,$ (16)
@@
    Fixing $\omega$, any problem whose PT estimator has higher variance possesses smaller $\omega_{j}$ and shrinks more toward $\tilde{Z}^{f}_{j}$; a smaller variance increases $\omega_{j}$ and makes the final estimator closer to $\hat{\theta}_{j}^{\mathrm{PT}}$.
@@
    Fixing all the problems, increasing $\omega$ has an overall effect of recovering $\hat{\theta}_{j}^{\mathrm{PT}}$ for all $j$ (full recovery when $\omega\to\infty$), and setting $\omega=0$ recovers $\tilde{Z}^{f}_{j}$.
Points (i) and (ii) establish the conceptual importance of $\omega$. If we could choose $\omega$ in an optimal way, that is,
@@
then the resulting estimator $\hat{\boldsymbol{\theta}}_{\omega^{*}}^{\mathrm{PAS}}$ would satisfy all our desiderata. While this construction is not feasible since the compound risk function in (9) depends on the unknown $\boldsymbol{\eta},\boldsymbol{\theta}$, we can make progress by pursuing a classical statistical idea: we can develop an unbiased estimate of the compound risk (Mallows, 1973; Stein, 1981; Efron, 2004) and then use it as a surrogate for tuning $\omega$.
@@
$\mathrm{CURE}\left(\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\omega}\right)$ $:=\frac{1}{m}\sum_{j=1}^{m}\Big{[}(2\omega_{j}-1)\tilde{\sigma}^{2}_{j}+2(1-\omega_{j})\tilde{\gamma}_{j}$
$+(1-\omega_{j})^{2}\big{(}\hat{\theta}_{j,\omega_{j}}^{\mathrm{PT}}-\tilde{Z}^{f}_{j}\big{)}^{2}\Big{]}.$
Both the formula and our nomenclature (“correlation-aware”) highlight the fact that we must account for the potentially non-zero covariance between shrinkage source $\hat{\theta}^{\text{PT}}_{j}$ and target $\tilde{Z}^{f}_{j}$, which can be explicitly written down as
$\tilde{\gamma}_{j}:=\mathrm{Cov}_{\eta_{j}}[\hat{\theta}_{j}^{\text{PT}},\tilde{Z}_{j}^{f}]=\lambda_{j}^{*}\mathrm{Var}_{\eta_{j}}[\tilde{Z}_{j}^{f}]=\frac{\gamma_{j}}{n_{j}+N_{j}}.$ (17)
@@
See Appendices B and F.1 for the proof and motivation. With Theorem 4.1 in hand, we now have a systematic strategy of picking $\omega$ by minimizing CURE, following the paradigm of tuning parameter selection via minimization of an unbiased risk estimate (as advocated by, e.g. Li (1985); Donoho & Johnstone (1995); Xie et al. (2012); Candès et al. (2013); Ignatiadis & Wager (2019); Ghosh et al. (2025)): [^11] The connection to EB is the following. Xie et al. (2012) and Tibshirani & Rosset (2019) explain that James-Stein-type estimators may be derived by tuning $\sigma_{\theta}^{2}$ (in Section 3) via minimization of Stein’s (1981) unbiased risk estimate (SURE).
@@
0:  $(X_{ij},Y_{ij})_{i=1}^{n_{j}}$, $(\tilde{X}_{ij})_{i=1}^{N_{j}}$, $\gamma_{j},\tau_{j},\sigma_{j}$ for $j\in[m]$, predictive model $f$
@@
5:    $\lambda_{j}^{*}=\texttt{{get\_pt\_param}}(\gamma_{j},\tau_{j},n_{j},N_{j})$
@@
10:  $\hat{\omega}=\texttt{{get\_shrink\_param}}((\hat{\theta}_{j}^{\mathrm{PT}})_{j=1}^{m},(\tilde{Z}^{f}_{j})_{j=1}^{m},(\tilde{\sigma}_{j}^{2})_{j=1}^{m})$
@@
Even though $\hat{\omega}$ does not admit a closed-form expression, the one-dimensional minimization can be efficiently carried out numerically (e.g., grid search). The final PAS estimator is:
@@
5 Asymptotic Optimality
In (18), we proposed selecting $\hat{\omega}$ by optimizing an unbiased surrogate of true risk. In this section, we justify this procedure theoretically. Our first result establishes that CURE approximates the true loss (whose expectation is the compound risk in (9)) uniformly in $\omega$ as we consider more and more problems.